OpenStudy (anonymous):

Simplify the following expression. √-144√-48 I don't understand how you came up with the answer..??

OpenStudy (anonymous):

√-144 = √144 * √-1 = 12i where i = √-1 √-48 = √16*√-3 = 4√3 i so √-144√-48 = 12i * 4√3i = 48√3* i^2 = -48√3